OpenStudy (anonymous):

the 4th term is (2x + 10),5th term is (4x-4) and the 6th term is (8x + 40). Then calculate for the following (i).the value of x (II) the first ten terms in (a) Geometric progression (b) Arithmetic progression

OpenStudy (amistre64):

do we assume it arith to start out?

OpenStudy (amistre64):

otherwise i dont see a reason why x can t be anything it wants to be

OpenStudy (amistre64):

a sequence need not have a defining function to be a sequence

OpenStudy (amistre64):

now, if it IS arith; then we simply can plug these into the general formula and see what comes out of it

OpenStudy (amistre64):

suppose the a1= (2x + 10), a2= (4x-4), a3= (8x + 40) and are part of an arith seq. then an = a1 + d(n-1) a2 = a1 + d(2-1) (4x-4) = (2x + 10) + 3d 2x - 14 = 3d (2x-14)/3 = d

OpenStudy (amistre64):

a3= a1 + d(n-1) (8x + 40) = (2x+10) + (2x-14)(2)/3 6x + 30 = (4x-28)/3 18x + 90 = 4x-28 14x = -118 x = -118/14 = -59/7

OpenStudy (anonymous):

good try